\underline{Esercizio 5}

Date le funzioni f(x) = |x^{2} − 3x + 2| e g(x) = 2x − 1

• Dire quanto vale f(g(x)).f(g(x)) = |(2x − 1)^{2} − 3(2x − 1) + 2| = |4x^{2} − 10x + 4| • Dire quanto vale g(f(x)), g(f(x)) = 2|x^{2} − 3x + 2| − 1,

• disegnare un grafico qualitativo di f(x), di g(x), di f(g(x)) e di g(f(x))

Matematica - slide Anna Torre anno accademico 2008-09